# Disentangling spin excitation continua in classical and quantum magnets using 2D nonlinear spectroscopy

### v1

@@ I Introduction
Detecting quantum spin liquid (QSL) states in experiments has remained a difficult feat given that there are no direct spectroscopic probes for its fractionalized excitationsWitczak-Krempa et al. (2014); Savary and Balents (2016); Zhou et al. (2017); Broholm et al. (2020). The dynamical spin correlations captured by inelastic neutron scattering (INS) offer a potential pathway, predicting a continuum of excitations attributed to the fractionalization of the spins in the QSL stateKnolle et al. (2014, 2015). However, alternative phenomena - such as frustrated magnets with large thermal fluctuations, large unit-cell magnetic orders, or impurity-induced damping of magnons - can similarly manifest broad and featureless INS signalsZhang et al. (2023). Additionally, a continuum of excitations is not exclusive to quantum states of matter; classical spin liquids, which may arise from materials with large spin-$S$ moments, can produce qualitatively similar excitation continuaSamarakoon et al. (2017, 2018); Zhang et al. (2023). In principle, the excitation continuum of many quantum spin liquids should exhibit a momentum-dependent lower edge of the continuum, but INS may not have the low energy resolution to observe such a feature. Hence, it is difficult to pinpoint the exact origins of the appearance of a spin excitation continuum with INS, a probe that only accesses the linear response regime.
@@
Two-dimensional coherent spectroscopy (2DCS) has emerged as a powerful tool to explore the nonlinear response in various systemsMukamel (1995); Woerner et al. (2013); Grishunin et al. (2023). Instead of measuring the response to one pulse field as in conventional pump-probe experiments, multiple pump fields are employed, inducing higher-order optical processes, including the generation of higher harmonics as well as the production of sum and difference frequencies. Both theoretical and experimental studies have demonstrated sharp signals in the nonlinear spectroscopic response of various elementary excitations in quantum materials otherwise inaccessible from linear probesParameswaran and Gopalakrishnan (2020); Mahmood et al. (2021); Nandkishore et al. (2021); Gerken et al. (2022); Barbalas et al. (2023); Hart and Nandkishore (2023); Katsumi et al. (2023); Salvador et al. (2024); Wan and Armitage (2019); Watanabe et al. (2024); Lu et al. (2017); Zhang et al. (2024a); Choi et al. (2020); Qiang et al. (2023); Brenig and Krupnitska (2024). These include the superconducting Higgs mode in NbNKatsumi et al. (2023), Josephson plasmons in layered superconductorsSalvador et al. (2024), spinons in the transverse field Ising chainWan and Armitage (2019); Watanabe et al. (2024), higher harmonic generation from interacting magnons in a canted antiferromagnetLu et al. (2017); Zhang et al. (2024a, b), and Majorana fermions in the Kitaev spin liquid (KSL)Choi et al. (2020); Qiang et al. (2023). Notably, 2DCS on the Kitaev honeycomb model has uncovered sharp diagonal signals separated by 2-flux or 4-flux vison gaps. This naturally leads to the questions of whether 2DCS can illuminate the origins of the spin excitation continua seen with INS, and whether it can be used as a tool to rule out less exotic phenomena behind these measured continua.
@@
In this work, we provide insights into these questions by studying the 2DCS of classical frustrated magnets with large Kitaev interactionsKitaev (2006). It is well known that simulations of the dynamical spin structure factor (DSSF) from the classical Kitaev model produce qualitatively similar results as its quantum counterpartSamarakoon et al. (2017, 2018); Zhang et al. (2023). The similarities arise due to the momentum and energy dependence of the spin correlations being largely dictated by the degenerate classical ground state manifold. The primary difference between the classical and quantum KSLs originates from the 2-flux vison gap seen in the quantum model, which may be on the order of 0.6 meV if the Kitaev exchange is about 6 meV as in $\alpha-$RuCl3Kim and Kee (2016). Such a small gap would be difficult to resolve in INS experiments. Here, we first consider the pure Kitaev model at zero temperature and compare and contrast the 2D nonlinear spectroscopic responses in the classical and quantum models. For the classical model, we compute the 2DCS spectra using molecular dynamics (MD) simulations in conjunction with classical Monte Carlo simulations. Somewhat surprisingly, we observe sharp horizontal, vertical, diagonal, and anti-diagonal line signals in the 2D frequency space. Such sharp line features are characteristic of the quantum KSL, as reported earlierChoi et al. (2020); Qiang et al. (2023), and demonstrate that even classical dynamics can produce sharp “nonrephasing” and “rephasing” line features. However, the classical model does not give rise to the two-flux and four-flux gaps responsible for separating and shifting the signals in the quantum case. Furthermore, we see finite responses in various spin polarization channels where signals would be theoretically forbidden in the quantum model. Moving to finite temperature, we observe that the diagonal and anti-diagonal signals are highly sensitive to thermal fluctuations, with a temperature of $T/|K|=0.001$ already enough to wash out these signals.
@@
Lastly, we consider a minimal but realistic $K\Gamma\Gamma^{\prime}$ modelRau et al. (2014); Sears et al. (2015); Johnson et al. (2015) for the candidate Kitaev material $\alpha$-RuCl3 Plumb et al. (2014); Sandilands et al. (2015); Banerjee et al. (2016, 2017); Do et al. (2017); Wang et al. (2017); Banerjee et al. (2018); Zhou et al. (2023). A recent experiment applying high magnetic fields oriented perpendicular to the honeycomb plane reported a finite region of stability for a possible QSL phaseZhou et al. (2023). Based on theoretical studies of the $K\Gamma\Gamma^{\prime}$ model, this phase corresponds to a quantum paramagnet Gohlke et al. (2020), whose status as a spin liquid phase remains to be explored. In the classical limit, this region of the phase diagram corresponds to various competing large unit-cell magnetic orders at zero temperature, which form a thermal ensemble at finite temperatureChern et al. (2020). Remarkably, the DSSF of this ensemble of orders also yields an excitation continuumZhang et al. (2023), even though its origins are completely distinct from those from the KSL phase. Despite sharing a qualitatively similar DSSF, we observe strikingly distinct 2DCS responses for the classical KSL versus the ensemble of magnetic orders. The classical KSL shows sharp but smooth lines, in complete contrast to the dense collection of discrete points arising from magnons in the ordered phases. We therefore propose 2DCS as a potentially powerful probe for the underlying excitations of the putative QSL phase in $\alpha$-RuCl3 under an out-of-plane magnetic field.
@@ II Model and Methods
To simulate the dynamics induced by the incident pulse fields, we first study the static Hamiltonian in Eq. 1 using classical Monte Carlo methods, in which we treat the spins as vectors $\mathbf{S}_{i}=(S_{i}^{x},S_{i}^{y},S_{i}^{z})$. We then time evolve the resulting spin configurations in the presence of the pulse fields using the Landau-Lifshitz-Gilbert (LLG) equations of motionLakshmanan (2011),
@@ III.1 Quantum Model
where the nonlinear susceptibilities can be computed using time-dependent perturbation theory Mukamel (1995). For the $B_{A}^{z}B_{B}^{z}$ pulse configuration, the second order susceptibility $\chi^{(2),z}_{zz}(t,\tau)$ is zeroChoi et al. (2020), and thus the third order susceptibilities fully dictate the behavior of $\mathbf{M}_{NL}$. $\chi^{(3),z}_{zzz}(t_{3},t_{2},t_{1})$ is computed from the four-point correlation functions $R_{zzz}^{(l),z}$, given by
@@
×$\times\langle{0}|\hat{\sigma}_{m}^{z}|{P}\rangle\langle{P}|\hat{\sigma}_{l}^{z}|{Q}\rangle\langle{Q}|\hat{\sigma}_{k}^{z}|{R}\rangle\langle{R}|\hat{\sigma}_{j}^{z}|{0}\rangle$ (8)
@@
where $j,k,l,m$ are site indices and $P,Q,R$ label the energy eigenstates. Here, $|{0}\rangle$ is the ground state representing the zero flux and zero matter sector. The action of the local spin operators $\hat{\sigma}_{j}^{x}$,$\hat{\sigma}_{j}^{y}$,$\hat{\sigma}_{j}^{z}$ on $|{0}\rangle$ represents the creation of two fluxes across the $x$, $y$, or $z$ bonds respectively connected to site $j$, while also creating a local excitation in the matter sector. An illustration of a possible $\chi_{zzz}^{(3),z}(t,0,\tau)$ pathway is shown in Fig. 2A, and is an example of how one could deduce the energies of the intermediate excited eigenstates in the flux sector. For instance, the $A$ pulse polarized in the $z$ direction excites a matter fermion and two fluxes across the $z$ bond. Then, the $B$ pulse interacts twice, first creating a four-flux configuration, then annihilating a flux pair and returning to the two-flux sector. A measurement in the $z$ direction finally returns the two-flux state to the vacuum, obeying the cyclic property of the trace in Eq. III.1. Notably, trying to measure in the $\hat{\sigma}^{x}$ and $\hat{\sigma}^{y}$ channels for the $\mathbf{B}_{A}\parallel\hat{z}$, $\mathbf{B}_{B}\parallel\hat{z}$ pulse configuration will not yield a finite signal as there lacks an operation that would return us to the flux-free ground state. This principle gives rise to the selection rules in the quantum model: given two incident pulses in the $z$ direction, the only finite component of the magnetization is in the $z$ direction. Furthermore, the same principles dictate that the second order susceptibility must be zero with this pulse configuration.
@@
Equipped with the energies of the intermediate eigenstates, we can roughly predict where the signals should appear in $\chi^{(3),z}_{zzz}(\omega_{t},0,\omega_{\tau})$, a sketch of which is shown in Fig. 2B. Note that this sketch is an illustration of the constraints on $\chi^{(3),z}_{zzz}(\omega_{t},0,\omega_{\tau})$ rather than an actual calculation of the matrix elements. The main features from each correlator are the diagonal and antidiagonal signals separated by the two-flux or four-flux gaps. Furthermore, the $R_{zzz}^{(2,3),z}$ and $R_{zzz}^{(4),z}$ signals are offset in $\omega_{t}$ and $\omega_{\tau}$. Additionally, as reported earlierChoi et al. (2020), the $\chi^{(3),z}_{zzz}(\omega_{t},\omega_{\tau},0)$ channel (obtained from Fourier transforming $\chi^{(3),z}_{zzz}(t,\tau,0)$) gives rise to horizontal and vertical signals with offsets in $\omega_{t}$ and $\omega_{\tau}$ by two-flux or four-flux gaps (not shown here). With these features, along with the selection rules described above in mind, we will compare the results from the quantum model Choi et al. (2020); Qiang et al. (2023) to our MD results in the following subsections.
@@
Figure 3: Dynamical spin structure factor (A,B) and two-dimensional coherent spectroscopy (C-N) for the pure Kitaev models with $K=-1$ and $K=1$. A and B are replotted from Zhang et al. (2023), and were performed at $T/|K|=0.001$. $|M_{NL}^{x}(\omega_{t},\omega_{\tau})|$ (C-L), $|M_{NL}^{y}(\omega_{t},\omega_{\tau})|$ (D-M), and $|M_{NL}^{z}(\omega_{t},\omega_{\tau})|$ (E-N) are shown for $T/|K|=0$ and $T/|K|=0.001$. All plots were normalized to their respective maximum intensities. Before taking the Fourier transform, a Gaussian filter of $e^{-\eta(t^{2}+\tau^{2})}$ was applied, with $\eta=10^{-6}$.
@@ III.2 Classical Model: Zero Temperature
We first simulate the classical Kitaev model at zero temperature in order to directly compare against the quantum model. For ease of reference, we plot the dynamical spin structure factors for the ferromagnetic and antiferromagnetic Kitaev model from Zhang et al. (2023) in Fig. 3A and B, which qualitatively reproduce the results of the quantum model except for the two-flux gapSamarakoon et al. (2017); Yoshitake et al. (2020). The 2DCS for $K=-1$ and $K=1$ for three measurement channels ($x$, $y$, and $z$) are shown in Fig. 3C-E and Fig. 3I-K, respectively. In both cases, we see sharp diagonal nonrephasing signals in each measurement channel, and faint but sharp antidiagonal rephasing signals in the $z$ measurement channel (in Figs. 3E and K). The sharp diagonal signal is reminiscent of the $R^{(1),z}_{zzz}$ contribution from Fig. 2B in the quantum model, apart from the 2-flux gap. The horizontal and vertical signals are comparable to the contributions from the $\chi^{(3),z}_{zzz}(\omega_{t},\omega_{\tau},0)$ susceptibility, where the $A$ pulse interacts twiceChoi et al. (2020).
Figure 4: A, Even and B, odd decomposition of the 2DCS for the pure $K=-1$ model at zero temperature. $\textbf{M}^{+}_{NL}$ ($\textbf{M}^{-}_{NL}$) refers to the nonlinear response with two pulses polarized in the $+\hat{\textbf{z}}$ ($-\hat{\textbf{z}}$) direction. Before taking the Fourier transform, a Gaussian filter of $e^{-\eta(t^{2}+\tau^{2})}$ was applied, with $\eta=10^{-6}$.
We can immediately notice a number of discriminating features between the classical and quantum limits of the model. Firstly, since there is a finite signal in each measurement channel, the classical model does not obey the flux selection rules present in the quantum model (similar in spirit to the linear response case, in which off-diagonal $\chi^{(1),\alpha}_{\beta}$ vanish in the quantum model but are finite in the classical model). Secondly, although the classical model is able to produce sharp horizontal, vertical, diagonal, and antidiagonal signals, there is no offset of the signals by two- or four-flux gaps, as seen in Fig. 2B. In other words, the classical model cannot distinguish between the diagonal $R^{(1),z}_{zzz}$ and shifted diagonal $R^{(4),z}_{zzz}$ contributions.
One can further decompose the nonlinear response into its even and odd components, $\textbf{M}_{NL}=\textbf{M}_{NL,\text{even}}+\textbf{M}_{NL,\text{odd}}$, where $\textbf{M}_{NL,\text{even(odd)}}$ consists of the even (odd) order responses. By performing the simulation with the two pulses co-polarized in either the ±$\pm\hat{\textbf{z}}$ directions, the even/odd-order responses can be separated into
@@
where $\textbf{M}_{NL}^{\pm}$ are the nonlinear magnetizations with both pulses polarized in the ±$\pm\hat{\textbf{z}}$ direction. We present the separation of the even- and odd-order responses for the pure Kitaev model at zero temperature in Fig. 4. As mentioned earlier, the $\chi^{(2),z}_{zz}(t,\tau)$ (and higher even-order) response should be exactly zero for the quantum model due to the flux selection rulesChoi et al. (2020); Qiang et al. (2023). However, Fig. 4A shows a finite signal in the even-order response, again consistent with the fact that the classical model does not obey these selection rules. On the other hand, we see an antidiagonal signal appear only in the odd response shown in Fig. 4B. At leading order, this signal is unique to the third order $R^{(2,3),z}_{zzz}$ “rephasing” response in the quantum model, meaning that even at the classical level, we observe the sharp line features at the correct order.

### v2

@@ I Introduction
Detecting quantum spin liquid (QSL) states in experiments has remained a difficult feat given that there are no direct spectroscopic probes for its fractionalized excitationswitczak-krempa_correlated_2014; savary_quantum_2016; zhou_quantum_2017; broholm_quantum_2020. The dynamical spin correlations captured by inelastic neutron scattering (INS) offer a potential pathway, predicting a continuum of excitations attributed to the fractionalization of the spins in the QSL stateknolle_dynamics_2014; knolle_dynamics_2015. However, alternative phenomena - such as frustrated magnets with large thermal fluctuations, large unit-cell magnetic orders, or impurity-induced damping of magnons - can similarly manifest broad and featureless INS signalszhang_spin_2023. Additionally, a continuum of excitations is not exclusive to quantum states of matter; classical spin liquids, which may arise from materials with large spin-$S$ moments, can produce qualitatively similar excitation continuasamarakoon_comprehensive_2017; samarakoon_classical_2018; zhang_spin_2023. In principle, the excitation continuum of many quantum spin liquids should exhibit a momentum-dependent lower edge of the continuum, but INS may not have the low energy resolution to observe such a feature. Hence, it is difficult to pinpoint the exact origins of the appearance of a spin excitation continuum with INS, a probe that only accesses the linear response regime.
@@
Two-dimensional coherent spectroscopy (2DCS) has emerged as a powerful tool to explore the nonlinear response in various systemsmukamel_principles_1995; woerner_ultrafast_2013; grishunin_two-dimensional_2023. Instead of measuring the response to one pulse field as in conventional pump-probe experiments, multiple pump fields are employed, inducing higher-order optical processes, including the generation of higher harmonics as well as the production of sum and difference frequencies. Both theoretical and experimental studies have demonstrated sharp signals in the nonlinear spectroscopic response of various elementary excitations in quantum materials otherwise inaccessible from linear probesparameswaran_asymptotically_2020; mahmood_observation_2021; nandkishore_spectroscopic_2021; gerken_unique_2022; barbalas_energy_2023; hart_extracting_2023; katsumi_revealing_2023; salvador_principles_2024; wan_resolving_2019; watanabe_exploring_2024; lu_coherent_2017; zhang_terahertz_2024; choi_theory_2020; qiang_probing_2023; brenig_finite_2024. These include the superconducting Higgs mode in NbNkatsumi_revealing_2023, Josephson plasmons in layered superconductorssalvador_principles_2024, spinons in the transverse field Ising chainwan_resolving_2019; sim_nonlinear_2023; sim_microscopic_2023; watanabe_exploring_2024, higher harmonic generation from interacting magnons in a canted antiferromagnetlu_coherent_2017; zhang_terahertz_2024; nelson_upconv_2024, and Majorana fermions in the Kitaev spin liquid (KSL)choi_theory_2020; qiang_probing_2023. Notably, 2DCS on the Kitaev honeycomb model has uncovered sharp diagonal signals separated by 2-flux or 4-flux vison gaps. This naturally leads to the questions of whether 2DCS can illuminate the origins of the spin excitation continua seen with INS, and whether it can be used as a tool to rule out less exotic phenomena behind these measured continua.
@@
In this work, we provide insights into these questions by studying the 2DCS of classical frustrated magnets with large Kitaev interactionskitaev_anyons_2006. It is well known that simulations of the dynamical spin structure factor (DSSF) from the classical Kitaev model produce qualitatively similar results as its quantum counterpartsamarakoon_comprehensive_2017; samarakoon_classical_2018; zhang_spin_2023; bai_magnetic_2019; zhang_dynamical_2019; franke_thermal_2022. The similarities arise due to the momentum and energy dependence of the spin correlations being largely dictated by the degenerate classical ground state manifold. The primary difference between the classical and quantum KSLs originates from the 2-flux vison gap seen in the quantum model, which may be on the order of 0.6 meV if the Kitaev exchange is about 6 meV as in $\alpha$-RuCl3kim_crystal_2016. Such a small gap would be difficult to resolve in INS experiments. Here, we first consider the pure Kitaev model at zero temperature and compare and contrast the 2D nonlinear spectroscopic responses in the classical and quantum models. For the classical model, we compute the 2DCS spectra using molecular dynamics (MD) simulations in conjunction with classical Monte Carlo simulations. Somewhat surprisingly, we observe sharp horizontal, vertical, diagonal, and anti-diagonal line signals in the 2D frequency space. Such sharp line features are characteristic of the quantum KSL, as reported earlierchoi_theory_2020; qiang_probing_2023, and demonstrate that even classical dynamics can produce sharp “nonrephasing” and “rephasing” line features. However, the classical model does not give rise to the two-flux and four-flux gaps responsible for separating and shifting the signals in the quantum case. Furthermore, we see finite responses in various spin polarization channels where signals would be theoretically forbidden in the quantum model. Moving to finite temperature, we observe that the diagonal and anti-diagonal signals are highly sensitive to thermal fluctuations, with a temperature of $T/|K|=0.001$ already enough to wash out these signals.
@@
Lastly, we consider a minimal but realistic $K\Gamma\Gamma^{\prime}$ modelrau_generic_2014; sears_magnetic_2015; johnson_monoclinic_2015 for the candidate Kitaev material $\alpha$-RuCl3 plumb__2014; sandilands_scattering_2015; banerjee_proximate_2016; banerjee_neutron_2017; do_majorana_2017; wang_magnetic_2017; banerjee_excitations_2018; zhou_possible_2023. A recent experiment applying high magnetic fields oriented perpendicular to the honeycomb plane reported a finite region of stability for a possible QSL phasezhou_possible_2023. Based on theoretical studies of the $K\Gamma\Gamma^{\prime}$ model, this phase corresponds to a quantum paramagnet gohlke_emergence_2020, whose status as a spin liquid phase remains to be explored. In the classical limit, this region of the phase diagram corresponds to various competing large unit-cell magnetic orders at zero temperature, which form a thermal ensemble at finite temperaturechern_magnetic_2020. Remarkably, the DSSF of this ensemble of orders also yields an excitation continuumzhang_spin_2023, even though its origins are completely distinct from those from the KSL phase. Despite sharing a qualitatively similar DSSF, we observe strikingly distinct 2DCS responses for the classical KSL versus the ensemble of magnetic orders. The classical KSL shows sharp but smooth lines, in complete contrast to the dense collection of discrete points arising from magnons in the ordered phases. We therefore propose 2DCS as a potentially powerful probe for the underlying excitations of the putative QSL phase in $\alpha$-RuCl3 under an out-of-plane magnetic field.
@@ II Model and Methods
To simulate the dynamics induced by the incident pulse fields, we first study the static Hamiltonian in Eq. 1 using classical Monte Carlo methods, in which we treat the spins as vectors $\mathbf{S}_{i}=(S_{i}^{x},S_{i}^{y},S_{i}^{z})$. We then time evolve the resulting spin configurations in the presence of the pulse fields using the Landau-Lifshitz-Gilbert (LLG) equations of motionlakshmanan_fascinating_2011-1,
@@ III.1 Quantum Model
where the nonlinear susceptibilities can be computed using time-dependent perturbation theory mukamel_principles_1995. For the $B_{A}^{z}B_{B}^{z}$ pulse configuration, the second order susceptibility $\chi^{(2),z}_{zz}(t,\tau)$ is zerochoi_theory_2020, and thus the third order susceptibilities fully dictate the behavior of $\mathbf{M}_{NL}$. $\chi^{(3),z}_{zzz}(t_{3},t_{2},t_{1})$ is computed from the four-point correlation functions $R_{zzz}^{(l),z}$, given by
@@
$\times\langle{0}|\hat{\sigma}_{m}^{z}|{P}\rangle\langle{P}|\hat{\sigma}_{l}^{z}|{Q}\rangle\langle{Q}|\hat{\sigma}_{k}^{z}|{R}\rangle\langle{R}|\hat{\sigma}_{j}^{z}|{0}\rangle$ (8)
@@
where $j,k,l,m$ are site indices and $P,Q,R$ label the energy eigenstates. Here, $|{0}\rangle$ is the ground state representing the zero flux and zero matter sector. The action of the local spin operators $\hat{\sigma}_{j}^{x}$$\hat{\sigma}_{j}^{y}$$\hat{\sigma}_{j}^{z}$ on $|{0}\rangle$ represents the creation of two fluxes across the $x$, $y$, or $z$ bonds respectively connected to site $j$, while also creating a local excitation in the matter sector. An illustration of a possible $\chi_{zzz}^{(3),z}(t,0,\tau)$ pathway is shown in Fig. 2A, and is an example of how one could deduce the energies of the intermediate excited eigenstates in the flux sector. For instance, the $A$ pulse polarized in the $z$ direction excites a matter fermion and two fluxes across the $z$ bond. Then, the $B$ pulse interacts twice, first creating a four-flux configuration, then annihilating a flux pair and returning to the two-flux sector. A measurement in the $z$ direction finally returns the two-flux state to the vacuum, obeying the cyclic property of the trace in Eq. III.1. Notably, trying to measure in the $\hat{\sigma}^{x}$ and $\hat{\sigma}^{y}$ channels for the $\mathbf{B}_{A}\parallel\hat{z}$, $\mathbf{B}_{B}\parallel\hat{z}$ pulse configuration will not yield a finite signal as there lacks an operation that would return us to the flux-free ground state. This principle gives rise to the selection rules in the quantum model: given two incident pulses in the $z$ direction, the only finite component of the magnetization is in the $z$ direction. Furthermore, the same principles dictate that the second order susceptibility must be zero with this pulse configuration.
@@
Equipped with the energies of the intermediate eigenstates, we can roughly predict where the signals should appear in $\chi^{(3),z}_{zzz}(\omega_{t},0,\omega_{\tau})$, a sketch of which is shown in Fig. 2B. Note that this sketch is an illustration of the constraints on $\chi^{(3),z}_{zzz}(\omega_{t},0,\omega_{\tau})$ rather than an actual calculation of the matrix elements. The main features from each correlator are the diagonal and antidiagonal signals separated by the two-flux or four-flux gaps. Furthermore, the $R_{zzz}^{(2,3),z}$ and $R_{zzz}^{(4),z}$ signals are offset in $\omega_{t}$ and $\omega_{\tau}$. Additionally, as reported earlierchoi_theory_2020, the $\chi^{(3),z}_{zzz}(\omega_{t},\omega_{\tau},0)$ channel (obtained from Fourier transforming $\chi^{(3),z}_{zzz}(t,\tau,0)$) gives rise to horizontal and vertical signals with offsets in $\omega_{t}$ and $\omega_{\tau}$ by two-flux or four-flux gaps (not shown here). With these features, along with the selection rules described above in mind, we will compare the results from the quantum model choi_theory_2020; qiang_probing_2023 to our MD results in the following subsections.
@@
Figure 3: Dynamical spin structure factor (A,B) and two-dimensional coherent spectroscopy (C-N) for the pure Kitaev models with K=1K=-1 and $K=1$. A and B are replotted from zhang_spin_2023, and were performed at $T/|K|=0.001$. $|M_{NL}^{x}(\omega_{t},\omega_{\tau})|$ (C-L), $|M_{NL}^{y}(\omega_{t},\omega_{\tau})|$ (D-M), and $|M_{NL}^{z}(\omega_{t},\omega_{\tau})|$ (E-N) are shown for $T/|K|=0$ and $T/|K|=0.001$. All plots were normalized to their respective maximum intensities. Before taking the Fourier transform, a Gaussian filter of $e^{-\eta(t^{2}+\tau^{2})}$ was applied, with $\eta=10^{-6}$.
@@ III.2 Classical Model: Zero Temperature
We first simulate the classical Kitaev model at zero temperature in order to directly compare against the quantum model. For ease of reference, we plot the dynamical spin structure factors for the ferromagnetic and antiferromagnetic Kitaev model from zhang_spin_2023 in Fig. 3A and B, which qualitatively reproduce the results of the quantum model except for the two-flux gapsamarakoon_comprehensive_2017; yoshitake_majorana-magnon_2020. The 2DCS for K=1K=-1 and $K=1$ for three measurement channels ($x$, $y$, and $z$) are shown in Fig. 3C-E and Fig. 3I-K, respectively. In both cases, we see sharp diagonal nonrephasing signals in each measurement channel, and faint but sharp antidiagonal rephasing signals in the $z$ measurement channel (in Figs. 3E and K). The sharp diagonal signal is reminiscent of the $R^{(1),z}_{zzz}$ contribution from Fig. 2B in the quantum model, apart from the 2-flux gap. The horizontal and vertical signals are comparable to the contributions from the $\chi^{(3),z}_{zzz}(\omega_{t},\omega_{\tau},0)$ susceptibility, where the $A$ pulse interacts twicechoi_theory_2020.
Figure 4: A, Even and B, odd decomposition of the 2DCS for the pure K=1K=-1 model at zero temperature. $\textbf{M}^{+}_{NL}$ ($\textbf{M}^{-}_{NL}$) refers to the nonlinear response with two pulses polarized in the $\hat{\textbf{z}}$ ($\hat{\textbf{z}}$) direction. Before taking the Fourier transform, a Gaussian filter of $e^{-\eta(t^{2}+\tau^{2})}$ was applied, with $\eta=10^{-6}$.
We can immediately notice a number of discriminating features between the classical and quantum limits of the model. Firstly, since there is a finite signal in each measurement channel, the classical model does not obey the flux selection rules present in the quantum model (similar in spirit to the linear response case, in which off-diagonal $\chi^{(1),\alpha}_{\beta}$ vanish in the quantum model but are finite in the classical model). Secondly, although the classical model is able to produce sharp horizontal, vertical, diagonal, and antidiagonal signals, there is no offset of the signals by two- or four-flux gaps, as seen in Fig. 2B. In other words, the classical model cannot distinguish between the diagonal $R^{(1),z}_{zzz}$ and shifted diagonal $R^{(4),z}_{zzz}$ contributions. Note though that both the flux selection rules and hard flux gap discussed here are unique to the exactly solvable Kitaev model. Upon including perturbing interactions, these properties no longer necessarily apply, resulting in more similar quantum and classical 2DCS responses.
One can further decompose the nonlinear response into its even and odd components, $\textbf{M}_{NL}=\textbf{M}_{NL,\text{even}}+\textbf{M}_{NL,\text{odd}}$, where $\textbf{M}_{NL,\text{even(odd)}}$ consists of the even (odd) order responses. By performing the simulation with the two pulses co-polarized in either the $\pm\hat{\textbf{z}}$ directions, the even/odd-order responses can be separated into
@@
where $\textbf{M}_{NL}^{\pm}$ are the nonlinear magnetizations with both pulses polarized in the $\pm\hat{\textbf{z}}$ direction. We present the separation of the even- and odd-order responses for the pure Kitaev model at zero temperature in Fig. 4. As mentioned earlier, the $\chi^{(2),z}_{zz}(t,\tau)$ (and higher even-order) response should be exactly zero for the quantum model due to the flux selection ruleschoi_theory_2020; qiang_probing_2023. However, Fig. 4A shows a finite signal in the even-order response, again consistent with the fact that the classical model does not obey these selection rules. On the other hand, we see an antidiagonal signal appear only in the odd response shown in Fig. 4B. At leading order, this signal is unique to the third order $R^{(2,3),z}_{zzz}$ “rephasing” response in the quantum model, meaning that even at the classical level, we observe the sharp line features at the correct order.